Find how much money needs to be deposited now into an account to obtain $1,000(future value) in 10 years if the interest rate is 4.5% per year compounded continuously.
Answer by VFBundy(438)

1000 = P * e^(0.045*10)

1000 = P * e^(0.45)

1000 = P * 1.568

P = 1000/1.568

P = $637.76